OpenStudy (anonymous):

Two friends are cleaning a pool. Roger can clean it in 6 hours working alone. If Latonya works alone, it will take her 5 hours. How long will it take them working together? Round to the nearest tenth

OpenStudy (anonymous):

After one hour, Roger has done 1/6 and Latonya has done 1/5 Those two combined give 1/6+1/5=11/30 So 30/11= 2.7 hours. After 2.7 hours, the cleaning will be done.

OpenStudy (anonymous):

okay thank you